# Improve Value Estimation of Q Function and Reshape Reward with Monte Carlo Tree Search

### v1

@@ III-B Improve Q value estimation in DDQN and reshape rewards with Monte Carlo Tree Search
Some former research[23] [4] has explored the idea of combinations of Q-learning and MCTS, but mostly in simple perfect information games and not applicable to imperfect information games. Our algorithm introduces a more complex MCTS variant with modifications such as a different expansion procedure and alternative backpropagation methods. We have also modify loss functions during agent training and developed an MCTS-based reward shaping structure.
After expanding a new state, the simulation continues until the game ends in the traditional MCTS. We limit each simulation to only one state expansion since this is more computationally efficient. Unlike standard MCTS, a single agent typically plays through the entire game via self-play, with any new state being expanded, evaluated, and backpropagated. However, due to the varying information available to different players in Uno, this approach is not feasible. The player in the new state may be different, and it’s impossible for the next player to pass information back to the previous one if the two players are not the same.
In our approach, if the simulation starts with player 1’s turn and transitions to player 2’s, we skip evaluating or backpropagating from all states of player 2. Player 2 continues with their own strategy until it’s player 1’s turn again. If the game ends, we backpropagate the results from the last state where it was player 1’s turn along the path to the start state. Our algorithm ensures that all states along the MCTS simulation path belong to the player who starts the simulation.
The action taken by the agent in expansion is the legal action with highest sum of Q-value and the Upper Confidence bounds applied to Trees (UCT):
@@
where $Q(s,a)$ is the Q-value of the legal action $a$ under a state $s$, and $UCT(s,a)$ can be defined as:
@@
where $N(s)$ is the total number of visits of state $s$, and $N(s,a)$ is the total number of uses of legal action $a$ at the state $s$, while $c_{puct}$ is a constant representing the exploration term. When a new state is discovered, if it is not an end state, the Q-value will be predicted for each legal action under that state through neural networks. Q-value is passed from the child to the parent during backpropagation and it is updated based on:
$Q_{new}(s,a)=\frac{Q_{old}(s,a)\cdot N(s,a)+Q(s^{\prime},a^{\prime})+r(s^{\prime})}{N(s,a)+1},$ (4)
where $r(s^{\prime})$ is the immediate reward after taking $a$ to next state $s^{\prime}$, always 0 unless the game ends. If $s^{\prime}$ is the end state, $Q(s^{\prime},a^{\prime})$ will be 0, $r(s^{\prime})$ will be 1. The whole procedure of our MCTS in Uno follows (Graphical representation shown in Figure 5):
@@
    Selection: This process starts from the root (start state) to find a state that has not been expanded. Record the player ID of the root node, $ID_{root}$. If the current state is during player $ID_{root}$’s turn, we select the action with the highest $(Q+UCT)$ based on Equation 3 for simulation. If it is another player’s turn, that player will choose an action based on their own policy until they transition to a state belonging to player $ID_{root}$.
@@
    Expansion: This step unfolds on the previously unexpanded player $ID_{root}$’s state, selecting the action with the highest $(Q+UCT)$ value. We try to find a child, which is also a state of $ID_{root}$, of this unexpanded state. If a game-ending state occurs during another player’s turn, we immediately proceed to step 3. If not, the new state of player $ID_{root}$ is found, and this state is initialized by Q-value prediction for every legal action, after which we move to step 3.
@@
    Backpropagation: This step involves updating the Q-values based on Equation 4 for all states of player $ID_{root}$ along the path from the newest found state or end state to the root. Finally, repeat the above three steps for a new round of simulation.
In sampling of standard DDQN, When the agent is at current state $s$, actions are selected via $\epsilon$-greedy: with probability $\epsilon$, a random action is taken, and with $1-\epsilon$, the action with the highest Q-value (predicted by neural networks) is chosen. After agent transitioning to the next state $s^{\prime}$ and receiving reward $r$, the tuple $(s,a,s^{\prime},r)$ is stored for further training. The sampling process in our algorithm is divided into two parts: MCTS simulation and interaction with the real environment. Assume the current state $s$ in the real environment, which is also the start state of MCTS, is simulated by MCTS. After the simulation, each legal action $a$ under state $s$ is assigned a corresponding Q-value, $Q_{m}(s,a)$, based on our MCTS rules. We then use an epsilon-greedy strategy to select the action based on $Q_{m}(s,a)$. Unlike Q-values $Q(s,a)$ are derived from a single neural network in DDQN, our $Q_{m}(s,a)$ are obtained from repeated MCTS simulations, reducing overestimations of Q values and making them more accurate. During the simulation, whenever an end state is reached and a reward $r_{k}$ (-1 or 1) is obtained, we accumulate and average these rewards:
$r_{m}=\frac{\sum_{k=1}^{n}r_{k}}{N_{s}}$ (5)
,where $n$ is the number of times the reward received at end state, $N_{s}$ is the number of simulation and $r_{m}$ is the total average reward from MCTS. After the agent chooses action $a$ based on $Q_{m}(s,a)$ , it transitions to the next state $s^{\prime}$ and receives a reward $r$ from real environment. We then add $r$ and $r_{m}$ as the total reward $r_{t}$ after agent takes action $a$. In standard environment, the agent typically receives reward of 0 after taking an action during the ongoing game, but because of rules MCTS, our agent can receive proper rewards at certain points, preventing long periods without positive feedback. We then keep tuple $(Q_{m}(s,a),s,a,s^{\prime},r_{t})$ as training data.
@@
2:     Initialize data structures of $Q(s,a)$, $N(s,a)$, $N(s)$, $V(s)$
3:     $Q(s,a)$: set of Q-values by taking action $a$ under state $s$
@@
15:     $a_{\text{best}}\leftarrow\epsilon-greedy(Q(s,a))$
@@
17:     return $(Q(s,a_{best}),r_{m},a_{\text{best}})$
@@
25:         $V(s),Q(s,a)\leftarrow\textsc{PREDICT}(s)$triangleright Predict Q value of every action under state $s$
@@
29:     $a\leftarrow\arg\max_{a}\left(Q(s,a)+\text{UCT}(s,a)\right)$
@@
31:     $q\leftarrow\textsc{SIMULATE}(s^{\prime})$
32:     UPDATE($s$, $q$, $a$, $r$)
@@
42:function UPDATE($s$, $q$, $a$, $r$)
@@
44:     Update $Q(s,a)$triangleright Update Q-values with formula based on Equation 4
@@
Figure 5: The starting player is player 1, so we only expand and backpropagate values for player 1’s states.
$L=L_{\text{ddqn}}+L_{\text{mcts}}$ (6)
, where $L_{\text{ddqn}}$ is the same as in Equation 1, except that the $r$ in Equation 2 is replaced with the total reward $r_{t}$, and $L_{\text{mcts}}$ is:
$L_{mcts}=(Q_{m}(s,a)-Q_{\text{estimator}}(s,a;\theta))^{2}$
$Q_{m}(s,a)$ is Q values from MCTS.
@@ IV Experiments and Evaluations
We conducted experiments and trained our algorithm, DDQN with MCTS, alongside three traditional algorithms—DDQN, DMC, and NFSP—with the same number of training episodes in two Uno environments: a three-player game and a four-player game, configurations commonly played by human players. The performance of DDQN with MCTS was then compared to the three traditional algorithms. To ensure a fair comparison of the performance and learning capabilities of the four algorithms, during training of every reinforcement learning agent, there was only one corresponding RL agent in each environment, while the others were random-playing agents. The evaluation and comparison are based on two main metrics: Total Average Rewards and Win Rate. Rewards were given to the agents at the end of each game by the environment, with the agent receiving either -1 or 1. The total average reward was the sum of rewards accumulated across all games divided by number of games. The win rate was calculated as the number of games won by an agent divided by the total number of games played. Given the high randomness of the Uno game, the reward data exhibits significant variance; hence, we decided to mainly focus on the overall trends in the data. To reduce the effect of variance of data, after every 1,000 training episodes, the agent was tested by playing 1,000 games with agents playing cards randomly, and the total average rewards were recorded and plotted in the training graph.
@@
In the evaluation, each algorithm relied solely on its estimator network (DMC, DDQN, DDQN with MCTS) or policy network (NFSP) to select actions. Although DDQN with MCTS used MCTS during the sampling process, for fair comparison, it only used the trained estimator network to select the action with the highest Q-value. The code was written in Python 3.10, and agents were trained on a single NVIDIA RTX 4080 GPU. All agents shared the same neural network architecture and hyperparameters, consisting of fully connected layers with sizes 240x64, 64x64, and 64x61. The batch size was set to 32, the learning rate to 0.00005, and the reward discount factor to 0.99. During the sampling process for DDQN with MCTS, each state was simulated 50 times using MCTS. A higher number of simulations could lead to more backpropagation updates and potentially more accurate Q-value predictions, but it would also increase the sampling time. We selected 50 simulations as it provides a balance between having rewards $r_{m}$ based on Equation 5 in some simulations and avoiding a substantial increase in sampling time.
@@
We trained the four algorithms separately until their performance began to converge. The raw training graph and comparisons between our algorithm and the other three traditional methods in 3-player and 4-player games are shown in Figure 6. DDQN with MCTS achieves the highest average total reward greater than -0.05 in the 3-player game and greater than -0.25 in the 4-player game, indicating a win rate close to 50% in the 3-player game against random-playing agents, and a win rate higher than 37.5% in the 4-player game. In general, DDQN with MCTS consistently outperformed the other algorithms in both environments. As the number of players increases, the training difficulty for the agent also grows. While other algorithms tend to perform progressively worse in environments with more players, DDQN with MCTS consistently maintains a stable rate of improvement and learning. More importantly, DDQN with MCTS improves very quickly in the early stages, whereas other algorithms only begin to improve slowly in the middle and later stages. From the raw graph, it is evident that the data exhibits a large variance due to the randomness inherent in the game. We focus more on the overall trends in the data when comparing these algorithms. Figure 7 shows a comparison of the mean and variance of the training data for the algorithms. In terms of the mean comparison, DDQN with MCTS achieves a total average reward that is 0.6 to 1.3 higher than the other algorithms in the 3-player game, meaning its win rate in tests against random-playing agents is 3% to 6.5% higher than that of the other algorithms. In the 4-player game, DDQN with MCTS achieves a total average reward that is 0.7 to 1.3 higher, corresponding to a win rate 3.5% to 6.5% higher than that of the other algorithms in tests against random-playing agents. We also tested DDQN with MCTS against the other algorithms in 10,000 games to evaluate its win rate when competing directly with them. Regardless of which algorithm it was tested against, DDQN with MCTS always achieved a higher win rate. In the 3-player game, its win rate was 4% to 16% higher than that of the other algorithms, and in the 4-player game, it was 5% to 11% higher, shown in Table III and Table IV.
@@
IV-A Test with Human Players and Knowledge learned by DDQN with MCTS
In order to provide human players with an opportunity to compete against our agents and to assess the performance of our algorithms in competition with human players, we have developed a fully functional graphical user interface. Our interface is implemented using the built-in Python module Tkinter [7]. We are also implementing online multiplayer functionality using the User Datagram Protocol (UDP) [10], allowing human players from different locations to remotely compete against our algorithms in 2-players, 3-players and 4-players environments, exampled GUI shown in Figure 8.
@@ IV-A Test with Human Players and Knowledge learned by DDQN with MCTS
We conducted a series of 100 games for each algorithm in one-on-one competitions with average human players. We observed DDQN with MCTS have surpassed human performance, achieving approximately a 54% win rate against average human players. In both the 3-player and 4-player games, the win rate of the agent is comparable to that of human players. When we tested the DDQN with MCTS algorithms against human players, we observed that these algorithms indeed learned some patterns of the game, which typically reflect strategies also exhibited by human players. For example, when the agents possess many cards along with a ’wild draw 4’, they tend to play it immediately. This strategy increases the number of cards the next player holds, making it more challenging for them to win the round. When the agents are down to only two cards, and one of them is a ’wild’ or ’wild draw 4’, they will save these for the final play. This is because, regardless of the target card’s color or number, the game rules allow a player with a wild card to play it directly, thus creating a guaranteed winning situation by holding onto this card until the end.

### v2

@@ III-B Improve Q value estimation in DDQN and reshape rewards with Monte Carlo Tree Search
Some former research[24] [4] has explored the idea of combinations of Q-learning and MCTS, but in simple perfect information games and not applicable to imperfect information games. Our algorithm introduces a more complex MCTS variant with modifications such as a different expansion procedure and alternative backpropagation methods. We have also modified loss functions during agent training based on MCTS returns and developed an MCTS-based reward shaping structure.
In the traditional MCTS, the simulation continues until the game ends after expanding a new state, and a single agent typically plays through the entire game via self-play, with any new state being expanded, evaluated, and backpropagated. However, due to the varying information available to different players in Uno, this approach is not feasible. The player in the new state may be different, and it’s impossible for the next player to pass information back to the previous one if the two players are not the same.
In our approach, We limit each simulation to only one step expansion since this is more computationally efficient. If the simulation starts with player 1’s turn and transitions to player 2’s, we skip evaluating or backpropagating from all states of player 2. Player 2 continues with their own strategy until it’s player 1’s turn again. If the game ends in simulation, we backpropagate the results from the end state directly to last state where it was player 1’s turn and along the path to the start state. Our algorithm ensures that all states along the MCTS simulation path belong to the player who starts the simulation.
The action taken by the agent in expansion is the legal action with highest sum of Q value and the Upper Confidence bounds applied to Trees (UCT):
@@
where $Q(s,a)$ is the Q value of the legal action $a$ under a state $s$, and $UCT(s,a)$ can be defined as:
@@
where $N(s)$ is the total number of visits of state $s$, and $N(s,a)$ is the total number of uses of legal action $a$ at the state $s$, while $c_{puct}$ is a constant representing the exploration term. When a new state is discovered, if it is not an end state, the Q value $Q(s,a)$ will be predicted for each legal action under that state through neural networks. Q value is passed from the child to the parent during backpropagation and it is updated based on:
$Q_{new}(s,a)=\frac{Q_{old}(s,a)\cdot N(s,a)+Q_{back}(s^{\prime},a^{\prime})}{N(s,a)+1},$ (4)
where
$Q_{back}(s^{\prime},a^{\prime})=\lambda\arg\max_{a^{\prime}}Q(s^{\prime},a^{\prime})+r(s^{\prime})$
$\lambda$ is the reward discount factor. $r(s^{\prime})$ is the immediate reward after taking $a$ to child state $s^{\prime}$, always 0 unless the game ends. If $s^{\prime}$ is the end state, $Q(s^{\prime},a^{\prime})$ will be 0, $r(s^{\prime})$ will be 1 or -1. The whole procedure of our MCTS in Uno follows (Graphical representation shown in Figure 5):
@@
    Selection: Record the player ID of the root node (start state), $ID_{root}$. This process starts from the root to find a state that has not been expanded. If a state is during player $ID_{root}$’s turn, the agent selects the action with the highest $(Q+UCT)$ based on Equation 3 for simulation. If it is another player’s turn, that player will choose an action based on their own policy until they transition to a state belonging to player $ID_{root}$.
@@
    Expansion: This step unfolds on the previously unexpanded player $ID_{root}$’s state, selecting the action with the highest $(Q+UCT)$ value. We try to find a child, which is also a state of $ID_{root}$, of this unexpanded state. If a game-ending state occurs during another player’s turn, we immediately proceed to step 3. If not, the new state of player $ID_{root}$ is found, and this state is initialized by Q value prediction for every legal action, after which we move to step 3.
@@
    Backpropagation: This step involves updating the Q values based on Equation 4 for all states of player $ID_{root}$ along the path from the newest found state or end state to the root. Finally, repeat the above three steps for a new round of simulation.
In sampling of standard DDQN, when the agent is at current state $s$, actions are selected via $\epsilon$-greedy: with probability $\epsilon$, a random action is taken, and with $1-\epsilon$, the action with the highest Q value $Q(s,a)$ (predicted by neural networks) is chosen. After agent transitioning to the next state $s^{\prime}$ and receiving reward $r$, the tuple $(s,a,s^{\prime},r)$ is stored for further training. However, the sampling process in our algorithm is divided into two parts: MCTS simulation and interaction with the real environment. Assume current state $s$ in the real environment, which is also the start state of MCTS, is simulated by MCTS. After the simulation, each legal action $a$ under state $s$ is assigned a corresponding Q value, $Q_{m}(s,a)$, based on our MCTS rules. We then use an $\epsilon$-greedy strategy to select the action based on $Q_{m}(s,a)$. Unlike Q values $Q(s,a)$ are derived from a single neural network in DDQN, our $Q_{m}(s,a)$ are obtained from repeated MCTS simulations, reducing overestimations of Q values and making them more accurate by average many Q values backups. During the simulation, whenever an end state is reached and a reward $r_{k}$ (-1 or 1) is obtained, we accumulate and average these rewards:
$r_{m}=\frac{\sum_{k=1}^{n}r_{k}}{N_{s}},$ (5)
where $n$ is the number of times the reward received at end state, $N_{s}$ is the number of simulation and $r_{m}$ is the total average reward from MCTS. After the agent chooses action $a$ based on $Q_{m}(s,a)$ , it transitions to the next state $s^{\prime}$ and receives a reward $r$ from real environment. We then combine $r$ and $r_{m}$ as the total reward $r_{t}$ after agent takes action $a$. In standard environment, the agent typically receives reward $r$ of 0 after taking an action during the ongoing game, but because of rules MCTS, our agent can receive proper rewards at certain points, preventing long periods without positive feedback. We then keep tuple $(Q_{m}(s,a),s,a,s^{\prime},r_{t})$ as training data.
Figure 5: The starting player is player 1, so we only expand and backpropagate values of player 1’s states.
$L=L_{\text{ddqn}}+L_{\text{mcts}},$ (6)
where $L_{\text{ddqn}}$ is the same as in Equation 1, except that the $r$ in Equation 2 is replaced with the total reward $r_{t}$, and $L_{\text{mcts}}$ is:
$L_{mcts}=(Q_{m}(s,a)-Q_{\text{estimator}}(s,a;\theta))^{2}$
@@
2:     Initialize data structures of $Q_{m}(s,a)$, $N(s,a)$, $N(s)$, $V(s)$
3:     $Q_{m}(s,a)$: set of Q values by taking action $a$ under state $s$
@@
15:     $a_{\text{best}}\leftarrow\epsilon-greedy(Q_{m}(s,a))$
@@
17:     return $(Q_{m}(s,a_{best}),r_{m},a_{\text{best}})$
@@
25:         $V(s),Q_{m}(s,a)\leftarrow\textsc{PREDICT}(s)$triangleright Predict Q value of every action under state $s$
@@
29:     $a\leftarrow\arg\max_{a}\left(Q_{m}(s,a)+\text{UCT}(s,a)\right)$triangleright a in $V(s)$
@@
31:     $Q_{m}(s^{\prime},a^{\prime})\leftarrow\textsc{SIMULATE}(s^{\prime})$
32:     UPDATE($s$, $a$, $r$, $Q_{m}(s^{\prime},a^{\prime})$)
@@
42:function UPDATE($s$, $a$, $r$, $Q_{m}(s^{\prime},a^{\prime})$)
@@
44:     Update $Q_{m}(s,a)$triangleright Update Q values with formula based on Equation 4
@@ IV Experiments and Evaluations
We conducted experiments and trained our algorithm DDQN with MCTS, alongside three traditional algorithms—DDQN, DMC, and NFSP—with the same number of training episodes in two Uno environments: a three-player game and a four-player game, configurations commonly played by human players. The performance of DDQN with MCTS was then compared to the three traditional algorithms. To ensure a fair comparison of the performance and learning capabilities of the four algorithms, during training of every reinforcement learning agent, there was only one corresponding RL agent in each environment, while the others were random-playing agents. The evaluation and comparison are based on two main metrics: Total Average Rewards and Win Rate. Rewards were given to the agents at the end of each game by the environment, with the agent receiving either 1 or -1. The total average reward was the sum of rewards accumulated across all games divided by number of games. The win rate was calculated as the number of games won by an agent divided by the total number of games played. Given the high randomness of the Uno game, the reward data exhibits significant variance. To reduce the effect of variance of rewards, after every 1,000 training episodes, the agent was tested by playing 1,000 games with agents playing cards randomly, and the total average rewards were recorded and plotted in the training graph.
@@
In the evaluation, each algorithm relied solely on its estimator network (DMC, DDQN, DDQN with MCTS) or policy network (NFSP) to select actions. Although DDQN with MCTS used MCTS during the sampling process, for fair comparison, it only used the trained estimator network to select the action with the highest Q value. The code was written in Python 3.10, and agents were trained on a single NVIDIA RTX 4080 GPU. All agents shared the same neural network architecture and hyperparameters, consisting of fully connected layers with sizes 240x64, 64x64, and 64x61. The batch size was set to 32, the learning rate to 0.00005, and the reward discount factor to 0.99. During the sampling process for DDQN with MCTS, each state was simulated 50 times using MCTS. A higher number of simulations could lead to more backpropagation updates and potentially more accurate Q value predictions, but it would also increase the sampling time. We selected 50 simulations as it provides a balance between having rewards $r_{m}$ based on Equation 5 in some simulations and avoiding a substantial increase in sampling time.
@@
We trained the four algorithms separately until their performance began to converge. The raw training graphs and comparisons between our algorithm and the other three traditional methods in 3-player and 4-player games are shown in Figure 6. DDQN with MCTS achieves the highest average total reward greater than -0.05 in the 3-player game and greater than -0.25 in the 4-player game, indicating a win rate close to 50% in the 3-player game against random-playing agents, and a win rate higher than 37.5% in the 4-player game. In general, DDQN with MCTS consistently outperformed the other algorithms in both environments. As the number of players increases, the training difficulty for the agent also grows. While other algorithms tend to perform progressively worse in environments with more players, DDQN with MCTS consistently maintains a stable rate of improvement and learning. More importantly, DDQN with MCTS improves very quickly in the early stages, whereas other algorithms only begin to improve slowly in the middle and later stages. We also focus on the overall trends in the data when comparing these algorithms. Figure 7 shows a comparison of the mean and variance of the training data for the algorithms. In terms of the mean comparison, DDQN with MCTS achieves a total average reward that is 0.6 to 1.3 higher than the other algorithms in the 3-player game, meaning its win rate in tests against random-playing agents is 3% to 6.5% higher than that of the other algorithms. In the 4-player game, DDQN with MCTS achieves a total average reward that is 0.7 to 1.3 higher, corresponding to a win rate 3.5% to 6.5% higher than that of the other algorithms in tests against random-playing agents. We also tested DDQN with MCTS against the other algorithms in 10,000 games to evaluate its win rate when competing directly with them. Regardless of which algorithm it was tested against, DDQN with MCTS always achieved a higher win rate. In the 3-player game, its win rate was 4% to 16% higher than that of the other algorithms, and in the 4-player game, it was 5% to 11% higher, shown in Table III and Table IV.
@@
IV-A Test with Average Human Players and Knowledge learned by DDQN with MCTS
In order to assess the performance of our algorithms in competition with human players, we have developed a fully functional graphical user interface (GUI). Our interface is implemented using the built-in Python module Tkinter [8]. We are also implementing online multiplayer functionality using the User Datagram Protocol (UDP) [11], allowing human players from different locations to remotely compete against our algorithms in 2-players, 3-players and 4-players environments, exampled GUI shown in Figure 8.
@@ IV-A Test with Average Human Players and Knowledge learned by DDQN with MCTS
We conducted a series of 100 games where our algorithm competed one-on-one against ourselves (representing average human players). We observed DDQN with MCTS have surpassed human performance, achieving approximately a 54% win rate against average human players. In both the 3-player and 4-player games, the win rate of the agent is comparable to that of human players. We also observed that these algorithms indeed learned some patterns of the game, which typically reflect strategies also exhibited by human players. For example, when the agents have many cards along with a ’wild draw 4’, they tend to play it immediately. This strategy increases the number of cards the next player holds, making it more challenging for them to win the round. When the agents are down to only two cards, and one of them is a ’wild’ or ’wild draw 4’, they will save these for the final play. This is because, regardless of the target card’s color or number, the game rules allow a player with a wild card to play it directly, thus creating a guaranteed winning situation by holding onto this card until the end.